OpenStudy (anonymous):

4. 8 – (3 + b) = b – 9 (1 point)

OpenStudy (anonymous):

8 - 3 - b = b - 9 distribute parenthesis 5 - b = b - 9 combine like terms 0 = - 14 no solution